# Using dust to constrain dark matter models

### v1

@@ 4.3 Dust
These two models were the Rémy-Ruyer et al. (2014) gas-to-dust ratio and the Li et al. (2019) best-fit set of parameters. In short, the dust-to-metals ratio produces the least total dust in all simulations, and the best-fit model produces the most dust in all simulations. All prescriptions produce highly concentrated dust distributions, though the transition in the gas-to-dust and best-fit prescriptions is more extreme from high to low densities. Despite these differences in the spatial distribution between dust mass prescriptions, we have confirmed they don’t significantly change the Gini scores. Further comparison of the dust models is provided in Appendix A.
@@ 4.4 Gini Trends
The difference in the Gini evolution produced by our simulations is an encouraging sign that the dust around a galaxy contains spatial information that can be used to differentiate between dark matter models. The individual Gini scores appear to be governed more strongly by the halo evolution rather than the baryonic model, which has two primary implications. Firstly, the diffuse dust is more dependent on the amount of substructure than the baryonic model. This is impacted by masking the central galaxy, where most of the dust is concentrated. For example, the gas disk is much more extended in the 5% warm dark matter simulation in Figure 3(b) compared to the 10% cold dark matter simulation in Figure 3(a). By masking this region to avoid saturating the Gini signal, we are also masking the region with the greatest difference caused by our implemented baryonic model.
The second implication is that every halo in our simulation has a unique evolution, even when starting from the same initial conditions and random seed. This unique evolution makes direct extrapolation of a Gini score to the observations difficult. However, the relative difference in the simulations highlights a difference that can be leveraged in observations. The absolute values for the total dust mass and Gini scores will likely change as a function of the random initial seed. Still, we expect that a warm dark matter universe, with fewer average mergers, will produce lower Gini scores over a large sample. Any prediction for an average Gini score in dust observations would require many simulations and careful treatment of the comparison between the two; though not beyond the capabilities of modern telescopes, with JWST and ALMA observations able to achieve sub-kpc resolution at (e.g. Liu et al., 2024). This is beyond the scope of the current work but will be the focus of our future work.
@@ 5 Conclusions
Our simulations take identical random seeds to create the initial conditions. The parent dark matter only simulations are then used to identify Milky Way analogues, hosted by haloes with a mass of approximately M⊙. Once identified, we ensure the haloes are isolated, to make sure there isn’t another large halo within 8 Mpc. After identifying target haloes in both the cold and warm dark matter parent simulations, the zoom-in simulations were run with 5% and 10% supernova efficiencies. The resimulated Milky Way analogue galaxies all show general agreement with observed quantities of the real Milky Way (see Table 2).
@@
The different supernova efficiencies produce different dust masses; the fiducial cold dark matter simulation produces the most at M⊙, compared to M⊙ in the fiducial warm dark matter simulation. All dust masses are listed in Table 3.
@@
The majority of the dust is always contained within the inner 15kpc of the halo, evident in Figure 3. We see a generally higher dust mass with cold dark matter across much of the halo (Figure 3(c)). There are regions with obvious differences, which are largely generated by differences in satellite positioning and an extended galactic disk in the warm dark matter simulation.
@@
The dust surface density profiles seen in Figure 4 show good agreement with the Ménard et al. (2010) scaling relation, , at radii larger than 20kpc. Between 10 to 20kpc, most simulations match reasonably well with the Draine et al. (2014) dust observations of M31, though, below these scales, the fiducial cold dark matter simulation has a rising density profile, whereas both warm dark matter simulations and the 5% cold dark matter simulation have flat density profiles.
@@
We calculated the Gini coefficient of each dust distribution, masking the inner 15kpc to study the diffuse dust, in Figure 5. We compared the evolution of the dust concentration across all four simulations, finding that a warm dark matter universe produces lower concentrations of dust. The evolution of the dust concentration is impacted more so by the merger history than the supernova efficiency, indicating these sorts of spatial studies can be leveraged by observations of dust around Milky Way-type galaxies.
@@ Acknowledgements
The authors would like to thank the referee, Mark Lovell, for their valuable comments in improving this publication, as well as useful discussions with Alexander Knebe and Desika Narayanan. AU and RM also acknowledge UWA/ICRAR for supporting a research visit. This research was supported in part by the Australian Government through the Australian Research Council Centre of Excellence for Dark Matter Particle Physics (CDM, CE200100008), and in part by the Australian Research Council Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D, CE170100013). AU acknowledges the support of an Australian Government Research Training Program Scholarship, and DC acknowledges the support of an ARC Future Fellowship (FT220100841). This work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government, and from the Victorian Higher Education State Investment Fund (VHESIF) provided by the Victorian Government.
@@ Appendix A Different dust production results
The different dust mass prescriptions available in powderday are discussed here, though we note they don’t impact the results of our analysis. In addition to the dust-to-metals method presented in Section 2.4, we also ran the metallicity-dependant gas-to-dust ratio described in Rémy-Ruyer et al. (2014), as well as the best-fit combined gas-to-dust and dust-to-metals ratios from the machine learning model described in Li et al. (2019). This best-fit model produces the most dust for each simulation out of any of the tested prescriptions, while the dust-to-metals ratio produces the least. For the cold dark matter simulation with 10% supernova efficiency, the best-fit model produced a total dust mass of M⊙, the gas-to-dust ratio a total mass of M⊙, compared with M⊙ for the dust-to-metals method.
We show the surface density profiles for all dust prescriptions in Figure 6. These all have similar shapes, though slightly different amplitudes. The best-fit surface density profile, Figure 6(c), has the greatest peak amplitude in the centre of the halo but has the lowest surface density at larger radii. All dust mass prescriptions follow the Ménard et al. (2010) scaling at radii 15kpc, though the dust-to-metals method produces comparatively higher masses in low density environments and so is fit better with a higher amplitude scaling.
More extreme changes in surface density are also evident in the Gini comparisons shown in Figure 7. The dust-to-metals method has the smoothest transition from the inner to outer halo in surface density (Figure 6(a)) and the least concentrated Gini evolution (Figure 7(a)). Conversely, the best-fit model has a more extreme transition in its density profile (Figure 6(c)) and a more concentrated Gini evolution (Figure 7(c)). The gas-to-dust dust mass calculation method is quite similar to the best-fit model with slightly lower concentrations of dust. Despite the differences in total dust mass and surface density profiles, the difference between warm and cold dark is persistent in all models, confirming our results above.
@@ Appendix B Halo masking
Here, we show the masking method used to calculate the Gini score outlined in Section 2.5. After creating the 2D projected histograms of dust mass similar to Figure 3, we take the centre of each bin, and if the centre of the bin is less than 15kpc from the centre of the halo, the bin is excluded from the calculation. Bins are also excluded if the bin centre is beyond R200c. The inner 15kpc and R200 limits are in comoving units for the evolution analysis. Figure 8 shows these masked 2D projections of the halo, with the masked regions in dark green. The same masking is applied to all rotations we use to calculate the median Gini score and errors.

### v2

@@ 4.3 Dust
These two models were the Rémy-Ruyer et al. (2014) gas-to-dust ratio and the Li et al. (2019) best-fitting set of parameters. In short, the dust-to-metals ratio produces the least total dust in all simulations, and the best-fitting model produces the most dust in all simulations. All prescriptions produce highly concentrated dust distributions, though the transition in the gas-to-dust and best-fitting prescriptions is more extreme from high to low densities. Despite these differences in the spatial distribution between dust mass prescriptions, we have confirmed they do not significantly change the Gini scores. Further comparison of the dust models is provided in Appendix A.
@@ 4.4 Gini trends
The difference in the Gini evolution produced by our simulations is an encouraging sign that the dust around a galaxy contains spatial information that can be used to differentiate between dark matter models. The individual Gini scores appear to be governed more strongly by the halo evolution rather than the baryonic model, which has two primary implications. First, the diffuse dust is more dependent on the amount of substructure than the baryonic model. This is impacted by masking the central galaxy, where most of the dust is concentrated. For example, the gas disc is much more extended in the 5 per cent WDM simulation in Fig. 3(b) compared to the 10 per cent CDM simulation in Fig. 3(a). By masking this region to avoid saturating the Gini signal, we are also masking the region with the greatest difference caused by our implemented baryonic model.
The second implication is that every halo in our simulation has a unique evolution, even when starting from the same initial conditions and random seed. This unique evolution makes direct extrapolation of a Gini score to the observations difficult. However, the relative difference in the simulations highlights a difference that can be leveraged in observations. The absolute values for the total dust mass and Gini scores will likely change as a function of the random initial seed. Still, we expect that a WDM universe, with fewer average mergers, will produce lower Gini scores over a large sample. Any prediction for an average Gini score in dust observations would require many simulations and careful treatment of the comparison between the two; though not beyond the capabilities of modern telescopes, with JWST and ALMA observations able to achieve sub-kpc resolution at (e.g. Liu et al., 2024). This is beyond the scope of the current work but will be the focus of our future work.
@@ 5 Conclusions
Our simulations take identical random seeds to create the initial conditions. The parent dark matter only simulations are then used to identify Milky Way analogues, hosted by haloes with a mass of approximately M⊙. Once identified, we ensure the haloes are isolated, to make sure there is not another large halo within 8 Mpc. After identifying target haloes in both the cold and warm dark matter parent simulations, the zoom-in simulations were run with 5 per cent and 10 per cent supernova efficiencies. The resimulated Milky Way analogue galaxies all show general agreement with observed quantities of the real Milky Way (see Table 2).
@@
The different supernova efficiencies produce different dust masses; the fiducial CDM simulation produces the most at M⊙, compared to M⊙ in the fiducial WDM simulation. All dust masses are listed in Table 3.
@@
The majority of the dust is always contained within the inner 15kpc of the halo, evident in Fig. 3. We see a generally higher dust mass with CDM across much of the halo (Fig. 3(c)). There are regions with obvious differences, which are largely generated by differences in satellite positioning and an extended galactic disc in the WDM simulation.
@@
The dust surface density profiles seen in Fig. 4 show good agreement with the Ménard et al. (2010) scaling relation, , at radii larger than 20kpc. Between 10 and 20kpc, most simulations match reasonably well with the Draine et al. (2014) dust observations of M31, though, below these scales, the fiducial CDM simulation has a rising density profile, whereas both WDM simulations and the 5 per cent CDM simulation have flat density profiles.
@@
We calculated the Gini coefficient of each dust distribution, masking the inner 15kpc to study the diffuse dust, in Fig. 5. We compared the evolution of the dust concentration across all four simulations, finding that a WDM universe produces lower concentrations of dust. The evolution of the dust concentration is impacted more so by the merger history than the supernova efficiency, indicating these sorts of spatial studies can be leveraged by observations of dust around Milky Way-type galaxies.
@@ Acknowledgements
The authors would like to thank the referee, Mark Lo v ell, for their valuable comments in improving this publication, as well as useful discussions with Alexander Knebe and Desika Narayanan. AU and RM also acknowledge the University of Western Australia/ICRAR for supporting a research visit. This research was supported in part by the Australian Government through the Australian Research Council Centre of Excellence for Dark Matter Particle Physics (CDM, CE200100008), and in part by the Australian Research Council Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D, CE170100013). AU acknowledges the support of an Australian Government Research Training Program Scholarship, and DC acknowledges the support of an ARC Future Fellowship (FT220100841). This work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government, and from the Victorian Higher Education State Investment Fund (VHESIF) provided by the Victorian Government.
@@ Appendix A Different dust production results
The different dust mass prescriptions available in powderday are discussed here, though we note they don’t impact the results of our analysis. In addition to the dust-to-metals method presented in Section 2.4, we also ran the metallicity-dependant gas-to-dust ratio described in Rémy-Ruyer et al. (2014), as well as the best-fit combined gas-to-dust and dust-to-metals ratios from the machine learning model described in Li et al. (2019). This best-fitting model produces the most dust for each simulation out of any of the tested prescriptions, while the dust-to-metals ratio produces the least. For the CDM simulation with 10 per cent supernova efficiency, the best-fit model produced a total dust mass of M⊙, the gas-to-dust ratio a total mass of M⊙, compared with M⊙ for the dust-to-metals method.
We show the surface density profiles for all dust prescriptions in Fig. 6. These all have similar shapes, though slightly different amplitudes. The best-fit surface density profile, Fig. 6(c), has the greatest peak amplitude in the centre of the halo but has the lowest surface density at larger radii. All dust mass prescriptions follow the Ménard et al. (2010) scaling at radii 15kpc, though the dust-to-metals method produces comparatively higher masses in low density environments and so is fit better with a higher amplitude scaling.
More extreme changes in surface density are also evident in the Gini comparisons shown in Fig. 7. The dust-to-metals method has the smoothest transition from the inner to outer halo in surface density (Fig. 6(a)) and the least concentrated Gini evolution (Fig. 7(a)). Conversely, the best-fit model has a more extreme transition in its density profile (Fig. 6(c)) and a more concentrated Gini evolution (Fig. 7(c)). The gas-to-dust dust mass calculation method is quite similar to the best-fit model with slightly lower concentrations of dust. Despite the differences in total dust mass and surface density profiles, the difference between warm and cold dark is persistent in all models, confirming our results above.
@@ Appendix B Halo masking
Here, we show the masking method used to calculate the Gini score outlined in Section 2.5. After creating the 2D projected histograms of dust mass similar to Fig. 3, we take the centre of each bin, and if the centre of the bin is less than 15kpc from the centre of the halo, the bin is excluded from the calculation. Bins are also excluded if the bin centre is beyond R200c. The inner 15kpc and R200 limits are in comoving units for the evolution analysis. Fig. 8 shows these masked 2D projections of the halo, with the masked regions in dark green. The same masking is applied to all rotations we use to calculate the median Gini score and errors.